# Identifying efficient routes to laminarization: an optimization approach

### v1

@@ II.2 Locating and tracking the edge
The minimal seed for transition and the minimal seed for relaminarization both lie infinitesimally close to the edge of chaos. However, the optimization-based approach detailed above for locating these minimal seeds involves sweeping over the perturbation magnitude, $d_{0}$, and solving a constrained nonlinear, nonconvex optimization problem for each value. Therefore, getting a close approximation of the minimal seeds and the edge using the optimization-based approach alone can become computationally expensive, as it requires solving this optimization over a fine discretization of this parameter space near the edge.
Rather than attempting to solve this challenging optimization problem over an increasingly fine discretization of $d_{0}$, we instead use the optimization method to obtain an approximation of these minimal seeds and then refine this approximation by applying an iterative bisection method that is commonly used for locating the edge of chaos [16, 20, 19, 21]. The idea behind the bisection approach is that given two initial conditions with one on each side of the edge, any simple path between these points must cross the edge. Therefore, one can bisect the line between these two points, classify which side of the edge the midpoint is on, and then replace one of the points with the midpoint, always keeping one point on both sides of the edge. Repeating this iteratively produces two points of arbitrarily small distance apart that lie on opposite sides of the edge. In many studies of the edge of chaos, this sort of iterative bisection is applied between the laminar state and an arbitrary initial condition that produces a chaotic transient. In our case, we can use the method to obtain a refined approximation of the minimal seed by using the minimal seed identified from the optimization procedure as one of the initial points in the iterative bisection scheme. Essentially, this assumes that the optimization approach identifies direction of the optimal perturbation and the bisection approach is then used to refine the approximation of the magnitude of the minimal seed.
Once a close approximation of the edge has been obtained, the iterative bisection strategy can be used to track trajectories along the edge of chaos. Since the edge separates qualitatively different trajectories, it will naturally be unstable. However, trajectories can be integrated from the two initial conditions on either side of the edge for a suitable time period. When it is found that these trajectories diverge from one another beyond the desired tolerance, they can be re-initialized using the iterative bisection method to reduce the distance between them to within the desired tolerance. In the context of the minimal seed problems, analyzing the edge trajectory near the minimal seeds can yield insights into what underlying dynamical structures influence the transition and relaminarization behavior.
@@ III Low-order model of a sinusoidal shear flow
Figure 3: Energetic quantities over time for a typical trajectory of the MFE system. From top to bottom: viscous dissipation rate, $D$; power input rate, II; kinetic energy, $E$; disturbance kinetic energy, $E_{d}$. $D$ and II values are normalized by their laminar value, $D_{L}$. Chaotic transient is denoted by gray, and the trajectory on the chaotic attractor is shown in cyan.
@@
Figure 4: Input-Dissipation projection for the trajectory of the MFE model showing sustained turbulent behavior shown as a time-series in Fig. 3. A chaotic initial transient is denoted by gray, and the trajectory on the chaotic attractor is shown in cyan. The initial condition for this trajectory is indicated by the black marker.
@@ IV Results
For the minimal seed for transition and minimal seed for relaminarization problems described above, we solve the optimization problem in Eqs. 2 and 3, respectively, over a range of evenly spaced $d_{0}$ values using the multi-step optimization approach detailed in Sec. II.1. For the multi-step method, we use windows of $\Delta t=10$ between breakpoints. Due to the nonconvex nature of these problems, the initialization of the optimization procedure can be quite important, especially when solving over long time horizons. For this reason, the problems are solved at sequentially increasing values of the final time, $\tau$, increasing in increments of $50$ time units, with the optimal perturbation for one $\tau$ value used as the initial guess in the optimization problem for the next $\tau$ value at a given perturbation magnitude, $d_{0}$. Each of these problems is solved using the penalty method described in Eq. 7 to convert the constrained optimization problem of Eq. 2 to a sequence of unconstrained subproblems with a penalty-augmented objective. These unconstrained subproblems are solved using the L-BFGS optimizer in PyTorch, with the needed gradients computed by backpropagation through the forward solution of the governing ODE using the torchdiffeq library [39]. The minimization of each of these subproblems is deemed sufficient when
@@
As noted above, to get a good approximation of the optimal perturbation using the optimization approach, it is necessary to solve these problems over a time horizon, $\tau$, that is sufficiently long for a trajectory to exhibit chaotic fluctuations for the transition problem or to approach the laminar state for the relaminarization problem. However, solving these problems becomes increasingly challenging for long time horizons due to the increased sensitivity of the objective with respect to the initial conditions. In practice, we have found that using the multi-step penalty method makes the optimization routine much more robust over longer horizons compared to directly solving the original problem without the multi-step penalty method. In some cases, careful choices of initial guesses and tuning of optimization parameters, including the the penalty parameter schedule and window lengths between breakpoints can be critical to achieving the desired convergence.
@@ IV.1 Minimal seed for transition
In the minimal seed for transition problem, as discussed in Sec. II, we search for the smallest perturbation of the laminar state that yields a trajectory that does not immediately laminarize. Fig. 5 (a) shows the final distance from laminar $d_{L}(\tau)$ as a function of the initial perturbation magnitude, $d_{L}(0)$. From this, the critical $d_{0}$ value at which the transition occurs can be identified by the point at which a sharp increase in $d_{L}(\tau)$ occurs. This signature of transition becomes increasingly clear for longer optimization horizons, as the jump becomes sharper and sharper – this has also been seen in other studies of the minimal seed for transition [28].
@@
Figure 6 shows the behavior of the two trajectories on either side of the edge at the minimal seed for transition, in terms of their distance from laminar. These trajectories remain quite close for approximately 500 time units before beginning to clearly separate, while the trajectory on the edge converges to a periodic orbit on the edge. This periodic orbit on the edge has also been identified for this system in Refs. [19, 21]. Fig. 7 shows the same two trajectories in the input-dissipation projection. These two trajectories remain close to each other as the approach the periodic orbit on the edge. Upon reaching this periodic orbit, the trajectories separate, each exiting in opposite directions along the unstable manifold of the periodic orbit on the edge. The low-side trajectory immediately laminarizes, while the high-side trajectory exhibits chaotic transience, visiting the vicinity of the chaotic attractor before eventually laminarizing. These results agree with the findings of Kim and Moehlis [19], where it was shown that the periodic orbit on the edge controls the transition behavior.
@@
Figure 7: Input-Dissipation projection for two trajectories on either side of the basin boundary near the minimal seed for transition (laminarizing , turbulent ). The minimal seed for transition is indicated by the blue marker at $t=0$ in panel (b). Also shown are a trajectory on the chaotic attractor ( ) and the edge periodic orbit ( ). (b) Zoom-in of the full-view in (a) near the edge periodic orbit. (c) Trajectories up to $t=500$ (vertical line in Fig. 6). (d) Trajectories for $t>500$ showing separation near the periodic orbit on the edge.
In Table 1, the magnitude of the minimal seed for transition, $\mathbf{\tilde{a}}\equiv\mathbf{a}-\mathbf{a}_{\mathrm{ref}}$, is reported along with the magnitude of each component of the normalized perturbation vector, $\hat{\mathbf{a}}=\mathbf{\tilde{a}}/\|\mathbf{\tilde{a}}\|$ , where the reference point $\mathbf{a}_{\mathrm{ref}}$ is the laminar state for the transition problem, and a point on the turbulent attractor for the relaminarization problem. From this, we see the optimal perturbation for triggering transition lies primarily in the direction of the $a_{3}$ mode which represents streamwise vortices, or rolls. This finding is in agreement of the view of the nonlinear self-sustaining process in wall-bounded turbulent flows proposed by Waleffe and coworkers [38, 24, 40, 41], where the formation of streamwise vortices plays an important role. In particular, in this view, weak streamwise vortices lead to the formation of streaks in the streamwise velocity. A nonlinear instability of these streaks then leads to streak oscillations and breakdown, which in turn leads to the regeneration of streamwise vortices through nonlinear effects, restarting the process. This calculation of the minimal seed for transition using the severely truncated MFE model also indicates that a weak excitation of the roll mode is the most efficient means of triggering transition from the laminar flow.
@@ IV.2 Minimal seeds for relaminarization
We now apply the same methodology to computing the minimal seed for relaminarization. In the minimal seed for relaminarization problem, we search for the smallest perturbation of the turbulent attractor that yields a trajectory that immediately laminarizes without a chaotic transient. For this, we solve the optimization problem in Eq. 3 over a range of evenly spaced perturbation magnitudes, $d_{0}$, using the multi-step optimization approach. This optimization problem requires a choice of reference point $\mathbf{x}_{T}$, in the turbulent region around which to compute the optimal perturbation. For most of this section, we take this to be the point on the turbulent attractor that is approximately the closest to the to the laminar state. We note that the optimal perturbation for achieving relaminarization will depend on the chosen reference point that is being perturbed. Our choice of reference here is somewhat arbitrary and other points may be more suitable depending upon the application being considered. At the end of this section, we briefly report results from other reference points in the turbulent attractor and in the turbulent region of the state space. While the magnitude and direction of the optimal perturbation varies depending on the chosen reference point, we find that the trajectories originating from the optimal perturbation at the minimal seed for relaminarization always take a similar laminarization route, passing near to the periodic orbit on the edge, regardless of the chosen reference point.
@@
To verify whether these trajectories truly laminarize, we consider the trajectories from these optimized initial conditions over a longer horizon of $800$ time units. Figure 8 (b) shows the resulting final distances from laminar for the initial conditions optimized over a horizon of $\tau=250$, simulated for $800$ time units. We evaluate the extended trajectories both in terms of their value at the final time, $t=800$, as well as the maximum distance from laminar taken over the last $500$ time units of the extended trajectory. Evaluating the trajectory at the longer final time allows us to ensure that the trajectories eventually converge to the laminar state, while evaluating the maximum distance from laminar allows us to rule out trajectories with a chaotic transient which eventually laminarize. From the distances at $t=800$ shown in Fig. 8 (b), it can be seen that two trajectories with perturbation magnitudes smaller than the critical magnitude where the drop occurs approach the laminar state by $t=800$. However, from observing the maximum distance from laminar from these two points in Fig. 8 (b), it is clear that these trajectories undergo a chaotic transient before approaching the laminar state. Furthermore, these evaluations of the extended trajectories also reveal that the true critical perturbation magnitude is slightly larger than expected from the trajectory evaluations over the optimization horizon of $\tau=250$. That is, two of the trajectories that appeared to laminarize in Fig. 8 ($d_{0}=0.012$ and $d_{0}=0.013$) do not truly laminarize. From Fig. 8 (b), the critical perturbation magnitude from the turbulent reference point for relaminarization is $d_{T}(0)=0.014$.
@@
To improve this approximation of the minimal seed for relaminarization, we use the bisection approach described in Sec. II.2 to locate a point on the edge between the turbulent reference point and the minimal seed approximation from the optimization approach. The optimization approach gives a point on the laminarizing side of the edge at a distance of $d_{T}=0.014$ from the turbulent reference point. We use this point and the reference point on the turbulent attractor to initialize the bisection algorithm. We then bisect along the line between these points, always keeping an initial condition on either side of the edge – one laminarizing immediately and one with at least a chaotic transient. We iteratively apply the bisection method until we obtain initial conditions separated by a distance of less than $10^{-6}$ and track the edge from these initial conditions by repeating the bisection whenever the distance between the trajectories grows beyond $10^{-6}$.
@@
Figure 9 shows the behavior of the two trajectories on either side of the edge found from this bisection approach, separated initially by less than $10^{-6}$, in terms of the distance of each trajectory from the laminar state over time. Also shown in Fig. 9 for reference, are the distance from laminar over time for a trajectory on the chaotic attractor and for the periodic orbit on the edge. The input-dissipation projections for these trajectories are shown in Fig. 10 (a) and (b). From this, it can be seen that these trajectories remain close together until approximately $t=600$ (vertical line in Fig. 9) before separating. The separation of the trajectories occurs as they approach the vicinity of the periodic orbit on the edge, with each one departing in different directions along the unstable manifold of this periodic orbit. Meanwhile, the edge trajectory beginning near the minimal seed for relaminarization converges to this periodic orbit. To highlight that this periodic orbit controls the laminarization behavior, in Fig. 10 (c) and (d), we show the trajectories beginning from initial conditions within a distance $10^{-6}$ of the edge found by re-initializing the bisection algorithm with points on the trajectories in Fig. 9 and Fig. 10 (a) and (b) at $t=600$ (vertical line in Fig. 9 and circle markers in Fig. 10 (b)). Fig. 10 (c) shows these trajectories as they approach the edge periodic orbit and Fig. 10 (d) shows the separation of these trajectories as they pass the edge periodic orbit.
@@
Figure 10: Input-Dissipation projection for two trajectories on either side of the basin boundary near the minimal seed for relaminarization (laminarizing , turbulent ). The minimal seed for relaminarization is indicated by the blue marker in panel (a). Also shown are a trajectory on the chaotic attractor ( ) and the edge periodic orbit ( ). (b) Zoom-in of the full-view in (a) near the edge periodic orbit. (c) Trajectories from initial conditions within $10^{-6}$ of the edge found by bisecting between points on the trajectories in (b) (circular markers). (d) Continuation of the trajectories in (c) as they separate at the edge periodic orbit.
Figure 11: Minimal seeds for relaminarization and laminarizing trajectories from multiple reference points in the turbulent region. The black markers show the reference points. Colored markers show the minimal seeds for each reference point. Also shown are a trajectory on the turbulent attractor ( ) and a periodic orbit in the turbulent region ( ). Other colored curves show the laminarizing trajectory from each minimal seed. The inset shows a magnified view of the region around the turbulent attractor.
@@
The laminarizing trajectory shown in Figs. 9 and 10 originates from the initial condition closest to the turbulent attractor that laminarizes without a chaotic transient. By tracking the nearby edge trajectory, the methods presented here reveal the dynamical mechanism that governs the relaminarization process. Specifically, the convergence of the edge trajectory to the periodic orbit on the edge indicates that the stable manifold of this orbit separates laminarizing trajectories from transiently turbulent trajectories. Previous works [19, 21] have identified this stable manifold as the edge of chaos for the MFE system, but those studies focus on initial conditions near the laminar state. By applying the optimization-based approach to approximate the edge near the turbulent attractor, we have identified a laminarization pathway that originates from within turbulent region, where edge geometry becomes increasingly complex. This type of trajectory could be particularly valuable from a control perspective, as it provides a reference trajectory which laminarizes naturally. Therefore, in order to achieve laminarization from other points on the turbulent attractor, one strategy could be to actuate the system to steer the trajectory toward the minimal seed for relaminarization or another point on the laminarizing trajectory. We anticipate that such a strategy would be efficient in terms of control effort, as once this minimal seed trajectory is reached, the system will laminarize naturally without further actuation. In this way, the minimal seed for relaminarization could serve as an interpretable, physically-based waypoint or target in an optimal control scheme for drag-reduction.

### v2

@@ II.2 Locating and tracking the edge
The minimal seed for transition and the minimal seed for relaminarization both lie infinitesimally close to the edge of chaos. However, the optimization-based approach detailed above for locating these minimal seeds involves sweeping over the perturbation magnitude, $d_{0}$, and solving a constrained nonlinear, nonconvex optimization problem for each value. Therefore, getting a close approximation of the minimal seeds and the edge using the optimization-based approach alone can become computationally expensive, as it requires solving this optimization over a small discrete increment of this parameter near the edge.
Rather than attempting to solve this challenging optimization problem with an increasingly small increment of $d_{0}$, we instead use the optimization method to obtain an approximation of these minimal seeds and then refine this approximation of the edge by applying an iterative bisection method. Such bisection methods are commonly used for locating the edge of chaos [16, 20, 19, 21]. The idea behind the bisection approach is that given two initial conditions with one on each side of the edge, any simple path between these points must cross the edge. Therefore, one can bisect the line between these two points, classify which side of the edge the midpoint is on, and then replace one of the points with the midpoint, always keeping one point on each side of the edge. Repeating this iteratively produces two points of arbitrarily small distance apart that lie on opposite sides of the edge. In many studies of the edge of chaos, this sort of iterative bisection is applied between the laminar state and an arbitrary initial condition that produces a chaotic transient. In our case, we can use the method to obtain a refined approximation of the minimal seed by using the minimal seed identified from the optimization procedure as one of the initial points in the iterative bisection scheme. Essentially, this assumes that the optimization approach identifies direction of the optimal perturbation and the bisection approach is then used to refine the approximation of the magnitude of the minimal seed.
Once a close approximation of the edge has been obtained, the iterative bisection strategy can be used to track trajectories along the edge of chaos. Since the edge separates qualitatively different trajectories, it will naturally be unstable. However, trajectories can be integrated from the two initial conditions on either side of the edge for a suitable time period. When these trajectories diverge from one another beyond the desired tolerance, they can be re-initialized using the iterative bisection method to reduce the distance between them to within the desired tolerance. In the context of the minimal seed problems, analyzing the edge trajectory near the minimal seeds can yield insights into what underlying dynamical structures influence the transition and relaminarization behavior.
@@ III Low-order model of a sinusoidal shear flow
Figure 3: Energetic quantities over time for a typical trajectory of the MFE system. From top to bottom: viscous dissipation rate, $D$; power input rate, II; kinetic energy, $E$; disturbance kinetic energy, $E_{d}$. $D$ and II values are normalized by their laminar value, $D_{L}$. Chaotic transient is denoted by gray ( ), and the trajectory on the chaotic attractor is shown in cyan ( ).
@@
Figure 4: Input-Dissipation projection for the trajectory of the MFE model showing sustained turbulent behavior shown as a time-series in Fig. 3. A chaotic initial transient is denoted by gray ( ), and the trajectory on the chaotic attractor is shown in cyan ( ). The initial condition for this trajectory is indicated by the black marker.
@@ IV Results
For the minimal seed for transition and minimal seed for relaminarization problems described above, we solve the optimization problem in Eqs. 2 and 3, respectively, over a range of evenly spaced $d_{0}$ values using the multi-step optimization approach detailed in Sec. II.1. For the multi-step method, we use windows of $\Delta t=10$ between breakpoints. This value is chosen to be short relative to the Lyapunov timescale associated with the transiently chaotic behavior – approximately 25 time units for the MFE system at these parameter values. Due to the nonconvex nature of these problems, the initialization of the optimization procedure can be quite important, especially when solving over long time horizons. For this reason, the problems are solved at sequentially increasing values of the final time, $\tau$, increasing in increments of $50$ time units, with the optimal perturbation for one $\tau$ value used as the initial guess in the optimization problem for the next $\tau$ value at a given perturbation magnitude, $d_{0}$. Each of these problems is solved using the penalty method described in Eq. 7 to convert the constrained optimization problem of Eq. 2 to a sequence of unconstrained subproblems with a penalty-augmented objective. These unconstrained subproblems are solved using the L-BFGS optimizer in PyTorch, with the needed gradients computed by backpropagation through the forward solution of the governing ODE using the torchdiffeq library [41]. The minimization of each of these subproblems is deemed sufficient when
@@
To get a good approximation of the minimal seed using the optimization approach, it is necessary to solve these problems over a time horizon, $\tau$, that is sufficiently long for a trajectory to exhibit chaotic fluctuations for the transition problem or to approach the laminar state for the relaminarization problem. However, solving these problems becomes increasingly challenging for long time horizons due to the increased sensitivity of the objective with respect to the initial conditions. In practice, we have found that using the multi-step penalty method makes the optimization routine much more robust over longer horizons compared to directly solving the original problem without the multi-step penalty method. In some cases, careful choices of initial guesses and tuning of optimization parameters, including the the penalty parameter schedule and window lengths between breakpoints can be critical to achieving the desired convergence.
@@ IV.1 Minimal seed for transition
In the minimal seed for transition problem, as discussed in Sec. II, we search for the closest point to the laminar state that yields a trajectory that does not immediately laminarize. Fig. 5 (a) shows the final distance from laminar $d_{L}(\tau)$ as a function of the initial perturbation magnitude, $d_{L}(0)$. From this, the critical $d_{0}$ value at which the transition occurs can be identified by the point at which a sharp increase in $d_{L}(\tau)$ occurs. This signature of transition becomes increasingly clear for longer optimization horizons, as the jump becomes sharper and sharper – this has also been seen in other studies of the minimal seed for transition [28].
@@
Figure 6 shows the behavior of the two trajectories on either side of the edge at the minimal seed for transition, in terms of their distance from laminar. These trajectories remain quite close for approximately 500 time units before beginning to clearly separate, while the trajectory on the edge converges to a periodic orbit on the edge. This periodic orbit on the edge has also been identified for this system in Refs. [19, 21]. Fig. 7 shows the same two trajectories in the input-dissipation projection. These two trajectories remain close to each other as they approach the periodic orbit on the edge. Upon reaching this periodic orbit, the trajectories separate, each exiting in opposite directions along the unstable manifold of the periodic orbit on the edge. The low-side trajectory immediately laminarizes, while the high-side trajectory exhibits chaotic transience, visiting the vicinity of the chaotic attractor before eventually laminarizing. These results agree with the findings of Kim and Moehlis [19], where it was shown that the periodic orbit on the edge controls the transition behavior. Note that our method only seeks to separate trajectories which laminarize immediately from those which exhibit transient chaos, but does not distinguish between the long-time behavior of trajectories which exhibit transient chaos (i.e., whether they eventually laminarize or collapse onto the chaotic attractor). In other words, we do not distinguish between a weak basin boundary and a strong basin boundary [22]. This is why the identified trajectory on the chaotic side of the edge exhibits only transient chaos before eventual laminarization.
(a)                         (b)
Figure 7: (a) Input-Dissipation projection for two trajectories on either side of the basin boundary near the minimal seed for transition (laminarizing , turbulent ). The minimal seed for transition is indicated by the blue marker at $t=0$ in panel (b). Also shown are a trajectory on the chaotic attractor ( ) and the edge periodic orbit ( ). (b) Zoom-in of the full-view in (a) near the edge periodic orbit.
In Table 1, the magnitude of the minimal seed for transition, $\mathbf{\tilde{a}}\equiv\mathbf{a}-\mathbf{a}_{\mathrm{ref}}$, is reported along with the magnitude of each component of the normalized perturbation vector, $\hat{\mathbf{a}}=\mathbf{\tilde{a}}/\|\mathbf{\tilde{a}}\|$ , where the reference point $\mathbf{a}_{\mathrm{ref}}$ is the laminar state for the transition problem, and a point on the turbulent attractor for the relaminarization problem. From this, we see the optimal perturbation for triggering transition lies primarily in the direction of the $a_{3}$ mode which represents streamwise vortices, or rolls. This finding is in agreement of the view of the nonlinear self-sustaining process in wall-bounded turbulent flows proposed by Waleffe and coworkers [40, 24, 42, 43], where the formation of streamwise vortices plays an important role. In particular, in this view, weak streamwise vortices lead to the formation of streaks in the streamwise velocity. A nonlinear instability of these streaks then leads to streak oscillations and breakdown, which in turn leads to the regeneration of streamwise vortices through nonlinear effects, restarting the process. This calculation of the minimal seed for transition using the MFE model also indicates that a weak excitation of the roll mode is the most efficient means of triggering transition from the laminar flow.
@@ IV.2 Minimal seeds for relaminarization
We now apply the same methodology to computing the minimal seed for relaminarization. In the minimal seed for relaminarization problem, we search for the closest point to a reference point on the turbulent attractor that yields a trajectory that immediately laminarizes without a chaotic transient. For this, we solve the optimization problem in Eq. 3 over a range of evenly spaced perturbation magnitudes, $d_{0}$, using the multi-step optimization approach. This optimization problem requires a choice of reference point $\mathbf{x}_{T}$, in the turbulent region around which to compute the optimal perturbation. For most of this section, we take this to be the point on the turbulent attractor that is approximately the closest to the to the laminar state. We note that the minimal seed for relaminarization will depend on the chosen reference point that is being perturbed. Our choice of reference here is somewhat arbitrary and other points may be more suitable depending upon the application being considered. Later in this section, we report results from other reference points in the turbulent attractor and in the turbulent region of the state space. While the magnitude and direction of the optimal perturbation varies depending on the chosen reference point, we find that the trajectories originating from the minimal seed for relaminarization always take a similar laminarization route, passing near to the periodic orbit on the edge, regardless of the chosen reference point.
@@
To verify whether these trajectories truly laminarize, we consider the trajectories from these optimized initial conditions over a longer horizon of $800$ time units. Figure 8 (b) shows the resulting final distances from laminar for the initial conditions optimized over a horizon of $\tau=250$, simulated for $800$ time units. We evaluate the extended trajectories both in terms of their value at the final time, $t=800$, as well as the maximum distance from laminar taken over the last $500$ time units of the extended trajectory. Evaluating the trajectory at the longer final time allows us to ensure that the trajectories eventually converge to the laminar state, while evaluating the maximum distance from laminar allows us to rule out trajectories with a chaotic transient which eventually laminarize. From the distances at $t=800$ shown in Fig. 8 (b), it can be seen that two trajectories with initial perturbation magnitudes smaller than the critical magnitude where the drop occurs approach the laminar state by $t=800$. However, from observing the maximum distance from laminar from these two points in Fig. 8 (b), it is clear that these trajectories undergo a chaotic transient before approaching the laminar state. Furthermore, these evaluations of the extended trajectories also reveal that the true critical perturbation magnitude is slightly larger than expected from the trajectory evaluations over the optimization horizon of $\tau=250$. That is, two of the trajectories that appeared to laminarize in Fig. 8 ($d_{0}=0.012$ and $d_{0}=0.013$) do not truly laminarize. From Fig. 8 (b), the critical perturbation magnitude from the turbulent reference point for relaminarization is $d_{T}(0)=0.014$.
To give a sense of how expensive the computations are, below we list the number of optimization iterations needed for the minimal seed for relaminarization results presented here. The total number of optimization iterations needed to produce Fig. 8 was $20,750$. Each of these optimization steps involves one forward solve of the ODE, backpropagation to obtain the gradient, and then multiple more forward solves of the ODE within a line search to determine the optimal step. There are 18 different $d_{T}(0)$ values considered in this plot, so, on average, this amounts to 1,144 iterations per $d_{T}(0)$ value considered. For just the $\tau=250$ case which is used in the later results, a total of $7,951$ optimization iterations were taken ($442$ per $d_{T}(0)$ value, on average). In general, for small $d_{T}(0)$ values, in which the initial condition is forced to lie near the chaotic attractor, the optimization is more challenging, taking as many as $1,561$ iterations for the $\tau=250$ case. On the other hand, for large $d_{T}(0)$ values, where an optimal can be found in the laminar basin of attraction, the optimization is significantly simpler, taking as few as $89$ iterations in some cases. We note that these values are not necessarily the most efficient or optimal values for obtaining these results, and the number of iterations required can vary significantly depending on the problem, the optimizer, and the hyperparameters of the optimization, but we report them to give a sense of the computational expense. We also note that the optimizations at different $d_{T}(0)$ values are easily parallelizable.
@@
To improve this approximation of the minimal seed for relaminarization, we use the bisection approach described in Sec. II.2 to locate a point on the edge between the turbulent reference point and the minimal seed approximation from the optimization approach. The optimization approach gives a point on the laminarizing side of the edge at a distance of $d_{T}=0.014$ from the turbulent reference point. We use this point and the reference point on the turbulent attractor to initialize the bisection algorithm. We then bisect along the line between these points, always keeping one initial condition on each side of the edge – one laminarizing immediately and one with at least a chaotic transient. We iteratively apply the bisection method until we obtain initial conditions separated by a distance of less than $10^{-6}$ and track the edge from these initial conditions by repeating the bisection whenever the distance between the trajectories grows beyond $10^{-6}$.
@@
Figure 9 shows the behavior of the two trajectories on either side of the edge found from this bisection approach, separated initially by less than $10^{-6}$, in terms of the distance of each trajectory from the laminar state over time. Also shown in Fig. 9 for reference, are the distance from laminar over time for a trajectory on the chaotic attractor and for the periodic orbit on the edge. The input-dissipation projection for these trajectories is shown in Fig. 10. From this, it can be seen that these trajectories remain close together until approximately $t=600$ (vertical line in Fig. 9) before separating. The separation of the trajectories occurs as they approach the vicinity of the periodic orbit on the edge, with each one departing in different directions along the unstable manifold of this periodic orbit. Meanwhile, the edge trajectory beginning near the minimal seed for relaminarization converges to this periodic orbit.
@@
Figure 10: Minimal seed for relaminarization: Input-Dissipation projection for two trajectories on either side of the basin boundary near the minimal seed for relaminarization (laminarizing , turbulent ). Also shown are a trajectory on the chaotic attractor ( ) and the edge periodic orbit ( ). Insets show a zoom-in near the edge periodic orbit (upper right) and the chaotic attractor (lower left).
@@
The laminarizing trajectory shown in Figs. 9 and 10 originates from the initial condition closest to the turbulent attractor that laminarizes without a chaotic transient. By tracking the nearby edge trajectory, the methods presented here reveal the dynamical mechanism that governs the relaminarization process. Specifically, the convergence of the edge trajectory to the periodic orbit on the edge indicates that the stable manifold of this orbit separates laminarizing trajectories from transiently turbulent trajectories. Previous works [19, 21] have identified this stable manifold as the edge of chaos for the MFE system, but those studies focus on initial conditions near the laminar state. By applying the optimization-based approach to approximate the edge near the turbulent attractor, we have identified a laminarization pathway that originates from within turbulent region, where edge geometry becomes increasingly complex. This type of trajectory could be particularly valuable from a control perspective, as it provides a reference trajectory which laminarizes naturally. This use case will be demonstrated in Sec. IV.2.1. We also note that the minimal seed for relaminarization lies in a region of the state space far from the laminar state and the edge periodic orbit and that this portion of the edge was not seen in tracking the edge from the minimal seed for transition (see Fig. 7). Our method has therefore enabled the discovery of a new part of the edge of chaos that is not explored by simply tracking the edge from the minimal seed for transition.
@@
Figure 11: Minimal seeds for relaminarization and laminarizing trajectories from multiple reference points in the turbulent region. The minimal seed trajectories vary in the vicinity of the chaotic attractor, but eventually follow a common path to laminar, passing near the periodic orbit on the edge. Also shown are a trajectory on the chaotic attractor ( ), a periodic orbit in the turbulent region, and the periodic orbit on the edge ( ). Other colored curves show the laminarizing trajectory from each minimal seed. Insets show magnified views of the region around the turbulent attractor and the edge periodic orbit.